### v4

# On the build-up of effective hyperuniformity from large globular colloidal aggregates

Abstract
A simple three-dimensional model of a fluid whose constituent particles interact via a short range attractive and long range repulsive potential is used to model the aggregation into large spherical-like clusters made up of hundreds of particles. The model can be thought of as a straightforward rendition of colloid flocculation into large spherical aggregates. We illustrate how temperature and particle density influence the cluster size distribution and affect inter- and intra-cluster dynamics. The system is shown to exhibit two well separated length and time scales, which can be tuned by the balance between repulsive and attractive forces. Interestingly, cluster aggregates at moderate/low temperatures approach a cluster glassy phase whereas cluster particles retain a local liquid-like structure. These states present a strong suppression of density fluctuations for a significant range of relatively large wavelengths, meeting the criterion of effective disordered hyperuniform materials as far as the intercluster structure is concerned.
I Introduction
Spontaneous pattern formation and self-limiting association can be deemed key phenomena both in biological and soft matter physicsStradner et al. (2004). In the former, colloidal suspensions are characteristic examples in which a variety of patterns have been identified in systems as disparate as amphiphilicsCiach et al. (2013), and proteins such as lysozimeLiu et al. (2010) or antibodiesYearley et al. (2014) (see Kovalchuk and coworkersKovalchuk et al. (2009) and references therein for a more comprehensive review of experimental examples). As a particularly interesting illustration of self-limiting aggregation, recently Sweatman and Lue Sweatman and Lue (2019) have argued that the formation of giant clusters due to the competition between attractive and repulsive interactions might well be behind the condensation of proteins/nucleic acids into membraneless organelles within the nuclei of eukaryotic cells Hirose et al. (2022). This is an alternative view to the “traditional” liquid-liquid phase separation picture that is common among the structural biology communityMurthy and Fawzi (2020).
As to pattern formation in soft matter systems, there is a growing interest on this subject, in particular for the possibilities that non-templated thermodynamically controlled nano-patterningSear et al. (1999); Boles et al. (2016); Barad et al. (2021) offers in the field of nanotechnology. In this connection, simple models that exhibit short range attraction and long range repulsion (SALR), Andelman et al. (1987); Seul and Andelman (1995) have shown to be able to reproduce the experimental behavior in great detail. In the pioneering work of Andelman and coworkers it was predicted by means of a mean field theory that, for a sufficiently large amplitude of repulsive interactions, modulated cluster phases will emergeArcher et al. (2007). The central role of attraction/repulsion competition in the build-up of globular bubble cluster, and ordered modulated phases or bicontinuous percolating clusters that span up to macroscopic sizes, has been known for more than three decadesSeul and Andelman (1995). The shape and size of the clusters and their spatial organization is directly determined by the balance between attractive and repulsive forces, as shown by the ground state calculations of Mossa et al.Mossa et al. (2004). Obviously, thermodynamic conditions (density and temperature/pressure) play a key role as well. Recently Liu and Xi Liu and Xi (2019) proposed a classification of SALR potentials in terms of the ratio between repulsive and attractive forces and particle size. Thus SALR potential are classified into type-I (range of the attraction is less than 20% of the particle size, and range of the repulsion is a fraction of the particle size), type-II (range of the attraction less than 20% of the particle size, range of repulsion is comparable or larger than particle size) and type-III (range of attraction beyond 20% of particle size and range of repulsion as in type II). In Ref. Liu and Xi (2019) it was shown that these three classes of interactions lead to specific features as to their phase behavior and cluster morphology.
The phase behavior of SALR systems has been dealt with extensively using mean field and density functional theory approaches Archer et al. (2007); Archer (2008); Archer et al. (2008); Bomont et al. (2012); Bomont and Costa (2012); Sweatman et al. (2014); Edelmann and Roth (2016); Pini and Parola (2017), effective field theoriesCiach (2008); Ciach et al. (2013); von Konigslow et al. (2013), and simulation approaches Archer and Wilding (2007); Klix et al. (2010); Godfrin et al. (2014); Zhuang and Charbonneau (2016); Zhuang et al. (2016). A comprehensive summary of relevant contributions in this connection can be found in recent review works Liu and Xi (2019); Ruiz-Franco and Zaccarelli (2021). It should be stressed that in all these instances, the long range repulsive component of the interaction is essential for the self-limitation of the aggregation process, preempting condensation/demixing transitions. The latter would drive the system towards spatially separate macroscopic phasesGodfrin et al. (2014) instead of the microstructuring induced by clustering. It is worth mentioning that other alternative approaches can explain self-limitation in aggregation processes as well. Such is the case, for instance, when the presence of specific interactions in mixtures limits the growth of the aggregate due the saturation of bonding sites by one of the componentsPalaia and Šarić (2022), in parallel with the chain termination in radical polymerization reactions.
Finally, as stated by Klix and coworkers Klix et al. (2010) SALR systems are intrinsically frustrated, and this also affects their dynamics, which is mostly controlled either by mesoscopic order or by metastable disorderKlix et al. (2010). These authors actually identify three dynamically arrested phases in their SALR model with long range electrostatic repulsions: a Wigner glass (moderate packing fraction/weak attractive interactions), a cluster glassy state (low packing fraction, strong attractive interactions), and a percolating gel phase (high packing fraction/strong attractive interactions). The possibility of the existence of a Wigner glass phase in disordered systems with long range electrostatic interactions was first postulated by Bose and Wilke using mode coupling theoryBosse and Wilke (1998). Dawson et al Dawson (2002) further analyzed in detail the possibility of dynamically arrested states in colloids stemming from either attractive or repulsive forces, whose ratio, as Klix et al. Klix et al. (2010) have shown, controls both the type of dynamics and topology of the glassy states.
Experimental evidence of colloidal glassy states is extensive and, for instance, numerous examples can be found in the review chapter by Weitz Weitz (2011). These glassy states are particularly interesting from the technological standpoint, since they provide an avenue to manufacture a wide range of solid-like materials starting from colloidal solutions, and as will be shown in this work, they offer a feasible alternative for the fabrication of disordered hyperuniform materials. These systems are exotic states, lying in between crystals and fluidsTorquato and Stillinger (2003); Torquato (2018), i.e. while being structurally disordered, they exhibit a hidden order that, in common with crystalline phases, suppresses large scale density fluctuations. Aside from its ubiquitous occurrence in physical and biological systems (cf. see Ref. Torquato (2018) for a comprehensive review of the multiple systems in which disordered hyperuniformity is present), hyperuniform materials have been shown to display particularly interesting opticalFlorescu et al. (2009a, b); Froufe-Pérez et al. (2017); Zhou et al. (2019); Milošević et al. (2019) and acoustic propertiesRomero-García et al. (2019); Chéron et al. (2022). Manufacturing hyperuniform materials from colloidal aggregates opens multiple paths to the fine tuning of their optical/acoustical properties, since in addition to the aggregate composition (and corresponding form factor of the constituent particles), the size and/or topology of the particulate aggregates can also be controlled by chemical and physical means.
In this article we will revisit a particularly simple model of colloidal SALR system that can yield all types of cluster phases, and ultimately will be shown to be capable of producing disordered effective hyperuniform states. The two dimensional version of this model was explored in detail two decades ago by Imperio and Reatto Imperio and Reatto (2004, 2006); Imperio et al. (2006); Imperio and Reatto (2007). The model, in which both attractive and repulsive interactions are simple exponentials, i.e. Kac interactionsKac (1959), was initially introduced in this context by Sear and coworkersSear et al. (1999) to account for the spontaneous patterning of quantum dots at the air-water interface. Later, Archer and coworkers Archer et al. (2007); Archer and Wilding (2007); Archer et al. (2008) used density functional theory and extensive Monte Carlo simulations in a three-dimensional version of the same model, in order to explore its phase behavior and map its corresponding phase diagram. Schwanzer and KahlSchwanzer and Kahl (2010) analyzed the competition between clustering and vapor liquid condensation, and the cluster/particle dynamicsSchwanzer et al. (2016) again in two dimensions. Interestingly, the same Kac-potential model was also used by Meyra and coworkers Meyra et al. (2012, 2015) to explain the formation of vegetation patterns in environments with limited resources. This illustrates how SALR effective interactions can provide a qualitative (and even quantitative) explanation for spontaneous patterning over various orders of magnitude in the spatial (and to some extent temporal) dimension. In addition to all these studies in the bulk, Bores et al. Lomba et al. (2014); Bores et al. (2015) performed simulations and theoretical studies on the influence of confinement into disordered media on the pattern formation for the same SALR model in two dimensions as well.
In the comprehensive work of Archer and Wilding Archer and Wilding (2007), in addition to the vapor-liquid transition, two first order transitions were identified depending on the packing fraction. Namely, a phase change at high dilution between a vapor and a fluid of liquid-like spherical clusters (globular cluster phase), and another transition between a liquid and fluid of spherical voids (bubble cluster phase). In this paper we will focus solely on the first instance, with interaction parameters and thermodynamic conditions tuned to yield relatively large clusters. Here, using large scale simulations we have analyzed the cluster size distributions, the intercluster and intracluster microscopic structure and dynamics illustrating their dependence on the thermodynamic conditions. In contrast with the two-dimensional model, in which as temperature is lowered both the intracluster and intercluster structure rapidly freeze into hexagonal arrangements, the three-dimensional system retains the liquid-like structure of the cluster droplets, and the freezing of intercluster structure is frustrated, with the system displaying features characteristics of a glass. In this work, we will exploit the long range of the net inter-cluster interactions and relatively large attractive interparticle interaction (a type-III SALR potential), to find the conditions in which clusters end-up in the vicinity of a dynamically arrested state, a cluster glassy phaseKlix et al. (2010). It will be shown that these states display a strong attenuation of density fluctuations for a significant region of small wavenumbers. The system effectively resembles a stealthy disordered hyperuniform materialTorquato and Stillinger (2003); Batten et al. (2008); Morse et al. (2023), i.e. due to the suppression of density fluctuations for a certain region of large wavelengths, the material becomes to some extent “invisible” to the corresponding radiation probes.
The rest of the paper is sketched as follows. In the next section we will describe the SALR model and the simulation conditions. An analysis of the onset of clustering is presented in Section III. The single cluster structure (cluster size distributions, density profiles, average life times) and dynamics will be reviewed in Section IV. Two-particle and intercluster correlations, and the analysis of hyperuniformity are commented upon in Section V. The article is closed with a summary of our most relevant conclusions and future prospects.
Figure 1: (Left) SALR interaction potential as described by Eq. (3). In the inset the Fourier transform of non-singular part of the potential, $\tilde{\phi}(Q)$, (black solid curve) is also represented. For comparison in dashed blue the corresponding transform of the untruncated potential is also plotted. In vertical lines we indicate the minima of both transforms which are the characteristic wave vectors for the truncated and untruncated interaction, namely $Q_{0}$ and $Q_{0}^{\infty}$. (Right) Average intercluster potential (solid curves) and cluster-particle potential (dashed curves), computed assuming an uniform average cluster density and spherical cluster radii.
Figure 2: Snapshot of a cluster configuration in the globular phase for $\rho\sigma^{3}=0.05$, $k_{B}T/\zeta=3$. Transparent spheres represent the excluded volume around the clusters (approx. twice the particle cluster diameter) resulting from the long range interactions. The effective cluster density inferred from the excluded volume is $\rho_{C}\lambda_{0}^{3}=0.82$, corresponding to a rather dense fluid.
II Model and methods
Figure 3: High temperature structure factor from our HNC calculations (lines) vs MD results (symbols). In these conditions the system is in the homogeneous phase displaying the initial formation of aggregates. A vertical line marks the position of the characteristic wave vector, $2\pi/\lambda_{0}$, The inset depicts the medium range structure factor, wherein intracluster correlations should appear. The values close to one imply absence of correlations, i.e. ideal gas behavior.
As mentioned in the Introduction the particles in our model will be interacting via a two-exponential Kac potential of the form used in Refs. Sear et al. (1999); Imperio and Reatto (2004); Archer et al. (2007); Archer and Wilding (2007); Bomont et al. (2012); Lomba et al. (2014) to which we have added a soft core as in Ref.Lomba et al. (2014) in order to facilitate the use of molecular dynamics, namely
$u^{SALR}(r)=\varepsilon\left[K_{r}e^{-\alpha_{r}r/\sigma}-K_{a}e^{-\alpha_{a}r/\sigma}+\left(\frac{\sigma}{r}\right)^{10}\right].$ (1)
Here $\sigma$ is a measure of the inner particle core diameter and will be used as unit length. Since this is a rather soft potential, the actual size, as estimated from the particle-particle distribution function is $\sigma_{eff}\approx 0.8\sigma$. As mentioned by Sear and coworkersSear et al. (1999), while the use of the attractive exponential can be justified as a more or less crude modelization of dispersive forces as in the classical works by Kac et al.Kac (1959); Kac et al. (1963), the repulsive exponential is used solely for the sake of computational simplicity. Thus, with $K_{a},K_{r}>0$, and $\alpha_{a}>\alpha_{r}>0$ one has the characteristics of a SALR potential with the presence of a maximum for a distance $d_{m}$, larger than that corresponding to the potential minimum, and whose analytical expression reads
$d_{m}=\frac{\sigma}{\alpha_{a}-\alpha_{r}}\log\frac{\alpha_{a}K_{a}}{\alpha_{r}K_{r}}.$ (2)
The effect of the inner repulsive core on the value of $d_{m}$ is negligible. $d_{m}$ is a very significant quantity that in the case of globular clusters approximately determines the cluster size, being a rough measure of its diameter. As will be explained below, due to the long range character of the interaction, the average intercluster distance is approximately twice as large, i.e. $\lambda_{m}=2d_{m}$. For computational convenience the potential has been truncated and shifted at $R_{c}=60\sigma$, i.e., we will be dealing with an interaction given by
$u(r)=\left\{\begin{array}[]{cc}u^{SALR}(r)-u^{SALR}(R_{c})&\mbox{if}\;r\leq R_{c}\\ 0&\mbox{if}\;r>R_{c}\end{array}\right.$ (3)
Setting $K_{r}=1$ and $K_{a}=2$, $\sigma=4\;$Å, $\varepsilon=0.1$ kcal/mol, $\alpha_{r}=0.1$ and $\alpha_{a}=0.25$ one gets the potential depicted in the left graph of Figure 1.
Alternatively, one can resort to the characteristic wave vector of the interaction (and its corresponding correlation length) which is the absolute minimum of the Fourier transform of its non-singular part, namely, the potential without the strong short range repulsion, $\phi(r)=u(r)-\varepsilon\left(\frac{\sigma}{r}\right)^{10}$. Note that for $r>2\sigma$ $u(r)$ and $\phi(r)$ are practically identical. Its connection with the structure factor and the build up of intermediate range order can easily be understood recalling that
$\lim_{Q\rightarrow 0}S(Q)=1+\rho\tilde{h}(Q)=\left[1-\rho\tilde{c}(Q)\right]^{-1}\approx\left[\chi_{0}^{-1}+\rho\tilde{\phi}(Q)/(k_{B}T)\right]^{-1}$ (4)
where the tilde denotes a Fourier transform, $\rho$ is the particle number density, $k_{B}$ is Boltzmann’s constant, and we have made use of the Ornstein-Zernike relationHansen and McDonald (2006) that connects the pair distribution function $g(r)=h(r)-1$ with the direct correlation function, $c(r)$. Additionally, $\chi_{0}$ represents the isothermal compressibility of a reference fluid of particles interacting with the repulsive short range part of the potential ($\propto r^{-10}$ in our case). This latter quantity stems from the random phase approximation (cf Ref. Archer et al. (2008)) which gives $c(r)\approx c_{0}(r)-u(r)/k_{B}T\approx c_{0}(r)-\phi(r)/k_{B}T$, being $c_{0}(r)$ the direct correlation function of the reference fluid. For our densities of interest $\chi_{0}\approx 1$. One sees that a minimum in the potential’s Fourier transform present at $Q_{0}$ implies a maximum at $S(Q_{0})$. If the denominator in the r.h,s. of (4) vanishes, the system reaches a Lifshitz point, corresponding to the onset of infinitely long ranged modulated phases. In our case one can easily determine the position of the minimum for the untruncated interaction (1), as
$Q_{0}^{\infty}$ $=$ $\sqrt{\frac{c\alpha_{a}^{2}-\alpha_{r}^{2}}{1-c}},\;\;\;c=\left(\frac{\alpha_{r}K_{r}}{\alpha_{a}K_{a}}\right)^{1/3}$
$\lambda_{0}^{\infty}$ $=$ $2\pi/Q_{0}^{\infty},$ (5)
being $\lambda_{0}^{\infty}$ being the correlation length for the system with untruncated interaction. In the inset of the left graph of Figure 1 one sees clearly the presence of a marked minimum in the untruncated potential at $Q_{0}^{\infty}\sigma=0.253$. Due to truncation this minimum is shifted in our case to $Q_{0}\sigma=0.337$. Remarkably, the correlation value estimated from $d_{m}$ (i.e. the potential maximum), namely, $0.2932\pi/\lambda_{m}=$ 0.293 is relatively close. This is simply due to fact that the lowest order approximation for the value of both quantities (in particular for $Q_{0}^{\infty}\sigma$) is $\alpha_{a}$, (here 0.25). This only applies to the specific functional form of our interaction and our choice of parameters ($\alpha_{a}K_{a}>>\alpha_{r}K_{r}$).
Now, in lamellar phases, $\lambda_{0}$ will determine the modulation distance, and in the bubble phase the average distance between the bubbles. In all cases the signature of this correlation length, $\lambda_{0}$, is the presence of a high intensity peak in the structure factor, S(Q), located at $Q_{0}$. With our choice of parameters $d_{m}=10.7\sigma$. The corresponding estimate from the characteristic wavevector $Q_{0}$ gives a distance $d_{0}=9.32\sigma$ relatively close to the former, The cluster diameter will lie somewhere between 9 and 11 molecular diameters. This will be further confirmed in the next Section when discussing the average cluster density profiles. We will consider systems at relatively low number densities in order to guarantee the formation of micellar-like phases. Two number densities will be studied, $\rho\sigma^{3}=0.025$, and 0.05. As expected from the difference between $Q_{0}^{\infty}$ and $Q_{0}$, the truncation of the SALR potential implies that cluster sizes will be somewhat smaller than those corresponding to an untruncated potential, also the corresponding correlation lengths will also be slightly smaller.
The fact that the $\lambda_{0}$ is approximately twice the position of the potential maximum is also easily understood in terms of average cluster-cluster interactions in the case of globular clusters. In this instance one can assume a uniform effective density, $\rho_{cl}^{eff}$, inside the spherical clusters of effective radius $R_{cl}^{eff}$. These two quantities can be considered as optimization parameters in a Reverse Monte Carlo procedureMcGreevy (2001), in which the interaction potential is computed from
[…]
with the objective function being the intercluster pair distribution. We have performed these calculations for the aforementioned densities. With the temperature defined as kBT/$\zeta$ in terms of the potential well depth, ($\zeta=0.068$ kcal/mol, see Figure 1), The cluster-cluster potentials have been evaluated for the lowest temperature considered, kBT/$\zeta=3$. The double volume integrals over the cluster volumes are carried out numerically using a Gauss-Legendre quadrature, with $u(r)$ given by Eqs. (1) and (3). With this procedure we end up with estimated effective cluster densities of $\rho_{cl}^{eff}\sigma^{3}=1.2$ ($\rho\sigma^{3}=0.05$) and 0.83 ($\rho\sigma^{3}=0.025$), with corresponding effective radii, $R_{cl}^{eff}/\sigma=4.5$, and 4, respectively. The results obtained from Eq. (6) are presented in the right graph of Figure 1. The dashed curves correspond to the cluster-particle interaction that is simply obtained by removing one of the effective cluster densities and one volume integration from Eq. (6). We see in the Figure that the net inter-cluster interaction is extremely repulsive, orders of magnitude more intense than the interparticle potential, and dies out at $\sim\lambda_{0}$. This explains why this quantity is closely connected with the intercluster separation, as reflected by the first peak of the cluster-cluster pair distribution that will be discussed in the next section. Our spherical clusters act like large repulsive spheres of an approximate exclusion diameter $\lambda_{0}$, whereas their actual diameter (the precise value depends on the definition as we will see) is $\approx\lambda_{0}/2$, i.e. the excluded volume is 23 times the average cluster volume. This is illustrated in Figure 2 where for a simulation snapshot the excluded volume is represented by a transparent sphere. The effective cluster density as estimated from the excluded volume is in this instance $\rho_{C}\lambda_{0}^{3}=0.82$, where $\rho_{C}$ is the number density of clusters. Note that this effective density corresponds to a rather dense fluid, in particular when compared with the very low net particle density ($\rho\sigma^{3}=0.05$).
Molecular dynamics simulations have been run specifically in the NVT ensemble using the LAMMPS packageThompson et al. (2021) for a total of 218000 particles of mass 100 amu. Distance and energy units for the simulation have been chosen following the LAMMPS package real units specification. The integration time step was set to 2 fs. Results will be presented in terms of reduced units, both for temperature (defined in terms of the well depth as indicated above), length (using $\sigma=4$Å  as unit length) and time, for which the reduced time unit is given by $\tau_{0}=(m\sigma^{2}/\zeta)^{1/2}=7.5$ps. The systems were equilibrated for 20 ns, of which the first 10 ns correspond to a slow cooling process following a linear temperature ramp starting from a fully disordered system at kBT/$\zeta=18$ down to the desired temperatures, namely, kBT/$\zeta$ = 3, 6, 9, 12, and 15 for $\rho\sigma^{3}=0.05$ and just kBT/$\zeta$ = 3, and 9 for $\rho\sigma^{3}=0.025$. It is worth stressing that shorter equilibration runs (5 ns) lead to very similar cluster size/radii distributions. This is an indication that the results presented correspond to equilibrium states. For much shorter equilibration runs and/or cooling times it is very easy to end up with bimodal or even multi-modal cluster size/radii distributions. Production runs were $\tau_{run}=20$ns long with configurations stored every 10 ps.
Figure 4: Average life time of the clusters, ($\tau_{av}$, as a function of temperature for $\rho\sigma^{3}=0.05$ (blue bars) and $\rho\sigma^{3}=0.025$ (red bars). The upper limit of the $y$-axis corresponds to the length of the simulation production run, $\tau_{run}$. Note that for $\rho\sigma^{3}=0.025$ the only temperatures calculated are $k_{B}T/\zeta=3,$ and 9.
Figure 5: a) Cluster size distribution dependence on temperature and global density b) Cluster radii of gyration distribution in terms of density and temperature. Average radii are indicated by vertical lines, with solid ones corresponding to $\rho\sigma^{3}=0.05$ and dash-dotted ones to $\rho\sigma^{3}=0.025$.
III Onset of clustering
As mentioned before, we have chosen potential parameters and thermodynamic conditions for the system to yield a globular cluster phase. We have first explored a series of temperatures for the two densities indicated above using the Hypernetted Chain integral equation Hansen and McDonald (2006). As shown by Archer and WildingArcher and Wilding (2007) this equation has a non-solution region which hides the cluster phase. This lack of solutions (in fact solutions have analytical continuation in complex spaceLomba and López-Martín (1995)) can be interpreted as a signature of a phase transition, in this instance a first order transition between the particle gas and a fluid of liquid-like spherical clusters. For a very closely related system, this first order transition was identified by Archer and WildingArcher and Wilding (2007) using both Density Functional Theory and extensive Grand Canonical Monte Carlo simulations exploiting multi-canonical reweighting aided by multi-histogram reweighting techniquesArcher and Wilding (2007). In our case we find that real solutions for the largest density disappear below $k_{B}T_{ns}^{HNC}/\zeta=14.9$, and $k_{B}T/\zeta=8$ for $\rho\sigma^{3}=0.025$. For $k_{B}T/\zeta=17.5$ and $\rho\sigma^{3}=0.05$ the HNC converges yielding a structure factor in excellent agreement with the simulation (cf. Figure 3). A visual inspection of simulation snapshots does not show evidence of clustering. Interestingly, in Figure 3 we first see that there is a significantly high prepeak (also known as intermediate range order peak Godfrin et al. (2014)) which reaches the value $S(Q_{max})\approx 2.7$ for $Q_{max}\approx 2\pi/\lambda_{0}$, whereas for $Q\sigma\gtrsim 1$ the structure factor corresponds to that of a dilute gas of uncorrelated particles (cf. the inset in Figure 3). Godfrin and coworkers propose the use of this precise value of the prepeak height as signature of the onset of the cluster phaseGodfrin et al. (2014). We will see in the next sections that lowering the temperature increases the height of the prepeak by orders of magnitude, i.e. $S(Q_{max})$ practically diverges for a $Q_{max}\sim Q_{0}\neq 0$, which suggests the existence of a non-zero Lifshitz pointBomont and Costa (2012). This means that the system becomes unstable with respect to a modulated inhomogeneous phase, whose modulation is given by $Q_{max}\approx 2\pi/\lambda_{0}$. The locus of Lifshitz points for varying densities yields the $\lambda$-line Archer and Wilding (2007), a term borrowed from the study of criticality in ionic fluidsStell (1995). It must be mentioned, that Bollinger and coworkersBollinger and Truskett (2016) suggested the use of an additional criterion to characterize the onset of the cluster phase, namely, a thermal correlation length with values within the interval $2\lesssim\xi_{T}/\sigma\lesssim 3$. This quantity can be estimated from a Lorentzian fit of the prepeak, which in our case yields $\xi_{T}/\sigma=5.6$, exceeding the limit suggested in Ref. Bollinger and Truskett (2016). However, one must bear in mind that Ref. Bollinger and Truskett (2016) is focused on Yukawa (screened Coulomb) interactions, and here we are dealing with bare exponentials. So far it has not been assessed to what extent the choice of the explicit form of the potential affects the reliability of this criterion. Until a complete analysis of the type performed in Ref. Bollinger and Truskett (2016) is carried out it seems reasonable to stick to Godfrin’s et al choice.
An illustration of a globular cluster configuration can be seen in Figure 2, where we have also represented the effective excluded volume of the clusters using transparent spheres. One can see that despite a relatively low concentration of clusters, the spheres representing the excluded volume are arranged in a rather dense packing, which will be reflected both in the microscopic structure and dynamics of our system.
IV Single-cluster structure and dynamics
In order to define the clusters we have used a geometric criterion, namely, all those particles whose separation is below $d_{cl}=1.3\sigma$ are considered as pertaining to the same cluster. This value is obtained from a preliminary run at the lowest temperature and $\rho\sigma^{3}=0.05$ for which the particle-particle distribution function, $g(r)$, was evaluated. We found that $d_{cl}=1.3\sigma$ roughly corresponds to the first non-zero minimum of $g(r)$, i.e. it is an estimate of the outer boundary of the first coordination shell. Once this clustering distance was set, we have run a GPU density based parallel scan (G-DBSCAN Andrade et al. (2013)) using an in-house code. Specific GPU optimizations were performed to calculate structure factors, pair distribution functions Levine et al. (2011); Sakharnykh (2015); Reuter and Köfinger (2019) and dynamic properties. Only clusters containing more than 4 particles have been considered for the analysis below.
A first observation from our cluster analysis is the fact that the structures are long-lived. In Figure 4 the average cluster life time, $\tau_{av}$ is displayed for the thermodynamics states under consideration. This quantity is defined as the average of the time a given cluster retains its identity (does not merge, split or dissolve, but it might incorporate/lose individual particles) in the course of the simulation. The top of the ordinate axis corresponds to the length of the production run, $\tau_{run}\approx 2700\tau_{0}$. For the highest density, up until $k_{B}T/\zeta=15$ the average lifetime exceeds half the length of the simulation, which is an indication of the stability of the structures. This is explained by the large depth of the cluster-particle potential (see right graph of Figure 1). At $k_{B}T_{c}/\zeta\approx 15$ the clusters are extremely short lived, this temperature being fairly close to the limit of real solutions of the Hypernetted Chain integral equation, which has been discussed above. For higher temperatures no significant clustering is found. Now, when the density decreases the average life time of the clusters decreases substantially. The real solution of the HNC disappears at $k_{B}T/\zeta=7.6$, but for $k_{b}T/\zeta=9$ the simulation results still display a significant degree of clustering, although much more short lived. Simulation runs for higher temperatures (results not included) do not show evidence of clusters satisfying our minimum size criterion. In Figures S1 in the Supplementary Information one can see snapshots of the evolution of the cluster formation as the system is cooled down from $k_{B}T/\zeta=18$ to $k_{B}T/\zeta=3$.
Figure 6: Temperature and density dependence of the average density profile of the clusters. Densities have been scaled with the effective diameter of the particles which is $\sigma_{eff}\approx 0.8\sigma$. Vertical lines denote the radii corresponding to the Gibbs dividing surface (cf. Eq. (8) in the text), with solid ones corresponding to $\rho\sigma^{3}=0.05$ and dash-dotted ones to $\rho\sigma^{3}=0.025$.
Figure 7: (Upper graph) Single particle mean square displacement dependence on temperature and density. (Lower graph) Density and temperature dependence of the center of mass mean square displacement for persistent clusters. In both graphs velocities and time are expressed in reduced units. In the insets the long time behavior is illustrated using log-log plots.
As to the cluster statistics, in Figure 5 we present the cluster size distribution for the two densities in question and two selected temperatures (left graph) and the corresponding cluster radii distributions depicted in the right graph. As estimate of the cluster radii we have used the value of the gyration radius, which is defined in the usual fashion IUP (2006) by
$R_{g}=\left<\sqrt{\frac{\sum_{i}m_{i}({\bf r}_{i}-{\bf R}_{com})^{2}}{\sum_{i}m_{i}}}\right>$ (7)
where the brackets indicate a ensemble average and ${\bf R}_{com}$ is the center of mass of the cluster whose radius of gyration is being computed. We see that for the largest density the size distribution peaks at around 450 particles and at 200 particles for the lowest. For all temperatures the clusters are liquid droplets, as will be shown below when analyzing the short time behavior of the mean square displacement.
Interestingly one can see that the shape of the cluster size distribution is hardly affected by temperature in the case of the largest density, whereas for low density it becomes much wider and much larger clusters appear as temperature rises. This, in combination with the short lifetime of the clusters seen in Figure 4, is an indication of a rapid merging and splitting of clusters taking place for the lower density/high T case. The large spread of the low density-high T size distribution is also reflected in the cluster radii distribution, which in this case is almost Gaussian-like. All remaining distributions are negatively skewed beta-distributions with a more or less wide hump at approximately one particle diameter below the maximum of $P(R_{g})$. Large clusters are in equilibrium with smaller ones, coalescing with them and letting particles escape their outer atmosphere to subsequently form smaller clusters. Merging of larger clusters is prevented by the large intercluster repulsion (cf. see right graph of Figure 1), and this is at the root of the beta-distribution shapes. For the low density/high temperature case thermal energy is sufficient to overcome the intercluster repulsion to a certain extent, and the equilibrium between destruction and creation of clusters leads to a Gaussian-like distribution.
This picture is complemented by an inspection of the cluster average density profiles, $\rho_{cl}(r)$, depicted in Figure 6. In this case, we have scaled the densities with $\sigma_{eff}=0.8\sigma$ in order to be able to compare the densities with those of standard fluids (e.g. for a Lennard-Jones fluid, liquid densities are usually assumed to be $\rho\sigma^{3}\gtrsim 0.4$). One first notices that despite the extremely low total particle densities, for $\rho\sigma^{3}=0.05$ the clusters display very high values of the effective density profile, even for high T. These values are well into the liquid domain. For the lower global density, the effective inner cluster density remains practically within the boundaries of the liquid state. In both cases there is a transition towards gas-like densities as the profile approaches the surface in a gradual fashion. Taking into account that the clusters are rather spherical, we have defined a Gibbs dividing surface in order to have a quantitative description of the separation between gas-like particles and liquid-like particles, as is customary in the description of gas-liquid interfaces Hansen and McDonald (2006). This surface places an ideal separation between the gas-like cluster region from its corresponding liquid-like counterpart. Here, for a spherical cluster this should be the surface of a sphere, whose radius (Gibbs dividing radius) $R_{Gibbs}$ is given by
[…]
Obviously, when one has a uniform density, $\rho_{cl}(r)=\rho_{cl}(0)$ with a sharp interface at $R_{Gibbs}$, both sides of Eq. (8) are identically zero.
We see that these Gibbs radii separating gas-like and liquid-like particles in the clusters (vertical lines in Figure 6) follow the same trends as the average radii of gyration (vertical lines in Figure 5), i.e. larger values for larger densities, and for the same density the radius increases with temperature. This latter trend reflects the larger spread of the cluster size distributions as temperature raises. The Gibbs radii are systematically larger ($\approx 0.5\sigma$) than the gyration radii. This is a consequence of the lesser weight of the “gas atmosphere” surrounding the quasi-spherical clusters in Eq. (7). It is worth noticing that the cluster density profiles reach vanishing values in the interval $4.5\sigma<r<5.5\sigma$, by which the average cluster spatial extent is $\approx\lambda_{0}/2$, in accordance with the estimated obtained from the potential’s characteristic wave vector (cf. Figure 1). The temperature and density dependence of both $R_{Gibbs}$ and $R_{g}$ found in this paper agrees with the finding of Schwanzer et al. Schwanzer et al. (2016) for much smaller clusters in two dimensions. It is important to stress that in our system further increases of density maintaining the temperature fixed would lead to an increase in the number and size of the clusters, that would eventually coalesce to form lamellar phases which finally mutate into a bubble phase. Other SALR interactions (or different potential parameters) at sufficiently low temperatures can yield a wide panoply of modulated structures, such as cluster crystals, gyroid phasesEdelmann and Roth (2016), crystals of bars, etc.
Figure 8: Vibrational frequency spectra derived from the particles velocity self-correlation at various temperatures and $\rho\sigma^{3}=0.05$. A detail of the low frequency region is presented in the inset in the lower right corner. The cluster vibrational frequency spectrum as computed from the velocity self-correlation of the persistent clusters centers of mass is depicted in the inset in the upper right corner.
Figure 9: Cluster-cluster pair distribution function and intra-cluster pair distribution function (inset) for the lowest temperature state $k_{B}T/\zeta=3$.
Now a few words concerning the single particle dynamics. In what follows, in order to make the results independent of specific details of the model, time and frequency will be reduced with the corresponding molecular time unit, $\tau_{0}$ and its inverse $\tau_{0}^{-1}$ respectively. Here we first look at the individual particle mean square displacement (m.s.d), plotted in the upper graph of Figure 7. It is apparent from the figure that there are clearly several separate regimes. First, there is a short time scale (under 2 $\tau_{0}$, i.e. for our system around 14ps) where the system is in a diffusive regime corresponding to intracluster movements (the mean square displacement is $\lesssim 16\sigma^{2}\approx<R_{g}>^{2}$). Then the system reaches a subdiffusive regime with a decreasing diffusion constant (slope of the curves). This reflects the finite size of the clusters. For much longer times (up to 2.7×$\times 10^{3}\tau_{0}$) the behavior of the mean square displacement is represented in the inset in log-log scale. The low density-high temperature case displays a fully linear behavior in the m.s.d. In this case the regime is consistently diffusive at all times, except for the short lived ballistic region at very short times, hardly visible in the graph. In contrast, the two systems with the largest density display a plateau after the initial diffusive and subdiffusive regimes, which is then followed by a long time slightly diffusive behavior, hardly visible in the case of low T and highest density. This long time behavior stems from the collective cluster movement which is shown in the lower graph of Figure 7. Here we have plotted the mean square displacement of the center of mass of persistent clusters, that is, those that preserve their identity during the length of the production run. At short times one can even appreciate the ballistic regime with a parabolic m.s.d. before the clusters begin to collide. We see that the long time behavior (inset in log-log scale) displays the same trends as the single particle mean square displacement. One can actually measure a tiny diffusivity for the clusters at higher density/low temperature, namely $D/(\sigma^{2}\zeta/m)^{1/2}=0.0003$. If this quantity is rescaled to take into account the average mass of the clusters we would still have a very small diffusivity, $D/(\sigma^{2}\zeta/M)^{1/2}\approx 0.006$. This is an indication that our system is approaching an arrested glassy state, and the dynamic slow down is not only due to the larger mass of the clusters. Taking into account that the source of this quasi-freezing of the cluster positions is mostly induced by the long range of the repulsive interactions (cf the right graph of Fig. 1) one is tempted to identify this state as a precursor of a Wigner glass of clustersRuiz-Franco and Zaccarelli (2021). However, strictly speaking the formation of Wigner crystals and glasses Duang and Guojun (2005) is due to extremely long ranged interactions such a Coulombic ones, in conjunction with higher packing fractions. Following the discussion of Klix and coworkers Klix et al. (2010) and bearing in mind the significant interparticle attraction and low particle packing fraction of our model, we are most likely somewhere near the dynamic transition towards a cluster glassy state. It is worth recalling that despite the very low particle density, as mentioned before, (cf Figure 2) due to the large intercluster repulsion the effective density as inferred from the cluster excluded volume lies in the range $\rho_{C}\lambda_{0}^{3}=0.74\sim 0.82$ for all cases studied here, corresponding to relatively dense fluid states.
These features are further confirmed by an analysis of the velocity self-correlation functions which are illustrated and commented upon in the Supplementary Information (cf. Figure S2). These quantities, once Fourier transformed provide the frequency spectra, which are plotted in Figure 8. We see first that the high values of $Z(0)$ confirm the liquid-like dynamics of the cluster constituent particles. For $k_{B}T/\zeta=3$ and $\rho\sigma^{3}=0.05$, cage effect vibrations are reflected as a tiny maximum (see lower inset of Figure 8) observed at $10^{12}(m\sigma^{2}/\zeta)^{1/2}\nu\approx 0.65$, whereas as T is increased a much wider maximum occurs at lower frequencies $10^{12}(m\sigma^{2}/\zeta)^{1/2}\nu\approx 0.25$. This maximum stems from the coupled particle-cluster dynamics, an effect also seen in the velocity self-correlation function depicted in Figure S2. In the upper inset we have the corresponding vibrational frequency spectrum derived from the persistent cluster velocity self-correlation function. Now the situation is completely different and the low frequency behavior corresponds to a that of a system approaching a glassy state, with very low diffusivities ($Z_{cl}(0)\longrightarrow 0$), in agreement with the long time behavior of the m.s.d. seen in Figure 7. Note that the frequency domain of the cluster dynamics is one order of magnitude smaller than that of single particles, reflecting the much slower dynamics of the clusters, a consequence of their large masses. The intensity of the peaks increases with temperature, as the kinetic energy increase reflects mostly in vibrations with larger amplitude around equilibrium positions, with the maxima shifting to slightly higher energies. This three maxima correspond to three minima (cages) in the cluster velocity self-correlation function (right graph in Figure S2), where also one can see that the cages become narrower and move to shorter times as temperature is increased. This effect is due to the higher vibrational energy of thermal origin present in the high T state, which for $\rho\sigma^{3}=0.05$ preserves some of the characteristics of a quasi-amorphous state. Lowering the density and/or further increasing the temperature above $k_{B}T/\zeta>15$ obviously ends up in the dissociation of the clusters and the complete disappearance of the vibrational structure of the spectrum.
Figure 10: Cluster-cluster structure factor for the lowest temperature state, $k_{B}T/\zeta=3$.. In inset the low-Q behavior is illustrated with using $Q^{2}$ as abscissa.
V Two-particle and intercluster structural correlations. Build-up of effective hyperuniformity
Figure 11: Density dependence of the total structure factor S(Q) (main graph) and the intracluster structure factor (upper right inset) for the system in stable cluster phases. In the lower left inset the low-Q behavior of the particle-particle structure factor is depicted.
We can now analyze two-particle correlations. To start with, in Figure 9 we present the cluster-cluster pair distribution function for the lowest temperature and the two densities under consideration. In the inset the corresponding particle-particle intracluster pair distribution functions are shown. Our previous discussion concerning the average intercluster potential (cf. right graph in Figure 1) explains that clusters are separated by the correlation length $\lambda_{0}\approx 20\sigma$. The smaller size of the clusters for low density implies the shift of the first peak of $g_{cl-cl}(r)$ to lower distances. The splitting of the second coordination shell for $\rho\sigma^{3}=0.05$ is a typical signature of the presence of an amorphous structurePan et al. (2011), although in our case there is a residual diffusivity. This feature is found in dense amorphous systems, but here the total number density is very low ($\rho\sigma^{3}=0.05$). However, we must bear in mind that the excluded volume of the clusters is much larger than what would correspond to their actual size, as previously discussed (cf. Figure 2). Recall that the effective cluster diameter (corresponding to the excluded volume) is approximately twice the spatial extent of the clusters. On the other hand, once the particle density is halved the splitting disappears and the system becomes more liquid-like, even if the cluster diffusivity remains relatively low ($D/(\sigma^{2}\zeta/m)^{1/2}=0.006$).
[…]
If we now move from real to Fourier space, in Figure 10 we have the cluster-cluster structure factor, $S_{cl-cl}(Q)$. A first feature is to be noted, the height of the first peak for both densities is beyond Hansen-Verlet’sHansen and Verlet (1969) freezing rule $S(Q_{max})\gtrsim 2.86$ for which a transition towards a crystal is to be expected. As discussed above, our system presents features of an amorphous solid, but with a somewhat peculiar structure factor. We see that $S(Q)\approx 0$, $\forall\;Q\lesssim 0.2\sigma^{-1}$. In fact, one finds that the hyperuniformity index $H=S(Q_{max})/S(0<Q\lesssim 0.2\sigma^{-1})\approx 10^{3}$ is considerably high. While this value is still away from the criterion for near hyperuniformity, $H\gtrsim 10^{4}$ proposed by Atkinson and coworkers Atkinson et al. (2016) in maximally random jammed systems, it represents a considerable attenuation of density fluctuations. Actually it meets the criterion of effective hyperuniformity, $H\gtrsim 10^{2}$ proposed by Chen and et al.Chen et al. (2018). As to the low-$Q$ limit, it is illustrated in the inset of Figure 10. We appreciate the a linear dependence with $Q^{2}$. Our system, being “effectively hyperuniform”, does not reach the limiting behavior $\lim_{Q\rightarrow 0}S_{cl-cl}(Q)\approx 0$, but we one gets closer as density increases (but this is hardly affected by temperature). This effect is most likely due to the increase of the effective repulsion, since the internal cluster density and average cluster size rise when the overall density is increased. The $Q^{2}$ dependence found in Figure 10 is consistent with a class I (effectively) hyperuniform material Torquato (2018). Local number fluctuations will further confirm this result as shown below.
Now a few words concerning the position of the first maximum of the structure factor, $Q_{max}$. We see that for $\rho\sigma^{3}=0.05$, $Q_{max}\sim 2\pi/\lambda_{0}$ whereas $Q_{max}(\rho\sigma^{3}=0.025)>2\pi/\lambda_{0}$. This reflects the smaller size of the clusters for the lower density (see the corresponding radii of gyration in Figure 5) and matches the shift to lower $r$ of the first maximum of $g_{cl-cl}$ in Figure 9. The quantity $\lambda_{0}$ is sort of a “ground state” estimate of the cluster size, as temperature rises or density decreases its accuracy is lower.
If we now look at the total structure factor, $S(Q)$ depicted in Figure 11, its most noticeable feature is the huge magnitude of the prepeak, which is the signature of a large degree of clustering. Obviously the prepeak positions correspond to that of the first maxima in the cluster-cluster structure factors discussed above. As to the issue of particle hyperuniformity, the situation is less clear. Even if the hyperuniformity index meets the criterion of effective hyperuniformity as defined above, in the inset one can appreciate that the small-Q values of S(Q) are far from small. The results for $\rho\sigma^{3}=0.05$ seem to approach zero as $Q\longrightarrow 0$, but our sample size does not allow for probing smaller $Q$-vectors. The scattering from small clusters and free particles seems to destroy hyperuniformity in a region where otherwise the cluster structure can be deemed as effectively hyperuniform (compare the region for $Q\sigma\lesssim 0.2$ in Figures 11 and 10).
Now, in the inset we can see the large wavenumber behavior of S(Q). The peak maxima in this region $Q_{max}\sigma\approx 8.1$ reflects interparticle correlations at $0.8\sigma$, which is precisely the effective particle diameter, $\sigma_{eff}$. This is in accordance with the maxima of $g(r)$ in the inset of Figure 9. Interestingly, $H=S(Q_{max})/S(Q_{min})=0.2\times 10^{2}$ which is not very far from the criterion of effective hyperuniformity. Note that both $Q_{max}$ and $Q_{min}$ correspond to the intracluster structure factor. A proper choice of the short range interparticle interactions might actually drive the system closer to exhibiting an additional wavenumber range where a considerable attenuation of radiation probes could be present.
As an additional assessment of the effective hyperuniform character of our systems, we have also analyzed the local density (or particle number) fluctuations, defined as
$\sigma^{2}_{N}(R)=<N^{2}>_{R}-<N>_{R}^{2}$ (9)
where $N$ refers to the number of particles and the subscript $R$ refers to the radius of a sampling spherical volume. This quantity is known to fulfillTorquato and Stillinger (2003) $\sigma^{2}_{N}(R)=<N>_{R}$ for a Poisson point pattern, which epitomizes a completely random disordered and uncorrelated system. In fact, in many disordered cases $\sigma^{2}_{N}(R)\propto R^{d}$, where $d$ is the dimensionality ($<N>_{R}\propto R^{d}$ in a uniform system). At the other end of the spectrum, one has point patterns known as hyperuniform that fulfill $\sigma^{2}_{N}(R)\propto R^{d-\alpha}$ with $\alpha>0$. Among these one finds regular patterns such as those of crystals and quasicrystals for which $\alpha=1$Torquato and Stillinger (2003). Other intermediate cases exist (e.g. with logarithmic scaling), with the term strong hyperuniformity reserved for those systems in which $\alpha\geq 1$ Oğuz et al. (2017). Back to our problem, in Figure 12 we present the corresponding analysis based on the configurations of the clusters’ center of mass. The number fluctuation is computed by averaging both the number of clusters, $N_{cl}$ and $N_{cl}^{2}$ for sample spheres of radius $R$ randomly placed in the simulation box. Note that particle number fluctuations are strongly affected by the periodic nature of our sample, since sampling over r-space is limited in accuracy by the Fourier space constraint that must be satisfied in the presence of periodic boundary conditions, namely, ${\bf Q}=(n_{x},n_{y},n_{z})2\pi/L$, where $n_{i}$ are integer numbers and $L$ is the side of the simulation box. For this reason, local number fluctuation can only be accurately computed for values well below half the simulation box size, otherwise effects of the periodic nature of the sample become visible. These limitations have been analyzed in detail by Wright in Ref. Wright (2017). Now, in the left graph of Figure 12, one observes that $\sigma^{2}_{N_{cl}}(R)\propto R^{2}$ except for the high temperature/low density case which exhibits a $R^{3}$ dependence. The curves exhibit strong oscillations, evenly distributed around the linear regression lines (dotted lines in the figure), and indicate that the system has strong spatial correlations consistent with a dense amorphous system. On the right graph the scaled ratio $\sigma^{2}_{N}(R)/(4\pi n_{cl}R^{3}/3)$ is displayed vs the radius of the sampling sphere, $R$. Here $n_{cl}=<N_{cl}>/L^{3}$ is the average density of clusters. A decreasing trend in the $R$-dependence of this ratio is a signature of hyperuniformity Torquato (2018). Both graphs in Figure 12 indicate that our systems (with the exception of the low density/high T case) display inter-cluster structures that can be deemed effective hyperuniform from the point of view of number (or local density) fluctuations. When the right graph of Figure 12 is compared with Figure 2 of Ref. Torquato (2018), we can clearly appreciate that our systems are somewhere in between ordered and disordered hyperuniform point configurations, as could be expected from a glassy-like system.
Figure 12: Left) Local particle number variation of cluster center of mass configurations vs the sample sphere radius squared, $R^{2}$. The linear dependence is an indication of a hyperuniform configuration. Deviation in linearity is apparent in the high T/low density case. Dotted lines represent linear regressions to the simulation data. Right) Local particle number variation of cluster center of mass configurations scaled with the corresponding random uniform value , $34\pi n_{cl}R^{3}/$ 3 vs the sample radius, $R$. A decaying trend indicates the presence of a disordered hyperuniform system. Again, for the high T/low density case the scaled number fluctuation tends to remain constant.
The same analysis can be performed on particle number fluctuations. These results are presented in similar graphs in Figure S3 of the Supplementary Information. Here the situation is somewhat inconclusive. On one hand, the decreasing trend of $\sigma^{2}_{N}(R)/(4\pi\rho R^{3}/3)$ vs $R$ might be an indication of the presence of a certain degree of hyperuniformity. On the other, the strong oscillations on the fluctuation itself (left graph of S3) prevent to reach a clear conclusion. Most likely, the traces of hyperuniformity result from the underlying effective hyperuniform structure of the clusters’ centers of mass.
One might conclude that from the point of view of cluster configurations, the systems studied (with the exception of the high temperature/low density one) can be cast into class I of effectively hyperuniform materials, since $\sigma^{2}_{N}(R)\sim R^{d-1}$. This corresponds to $\lim_{Q\rightarrow 0}S(Q)\propto Q^{\alpha}$ Torquato (2018) with $\alpha>1$. In our case the low-Q behavior matches $\alpha\sim 2$, even if the structure factors do not vanish completely in the zero wavenumber limit. As to the structure displayed by the particle fluid as a whole, no clear conclusion can be drawn at this stage.
VI Conclusions
In summary, we have presented a study of the structural and dynamic properties of a simple model of a type-III SALR self-associating fluid, focusing on the globular cluster phase that upon aggregation displays a cluster structure with effective hyperuniform disorder. Here we have studied both cluster size distributions, cluster density profiles, diffusion and frequency spectra as well as pair distributions (inter and intracluster) and structure factors. From the latter, we have found that the cluster phase for moderate/high cluster densities ($\rho_{C}\gtrsim 0.8$) approaches a cluster glassy state. In these conditions the cluster structure shows a considerable attenuation of density fluctuations for a range of wavenumbers below $0.2\sigma^{-1}$, meeting the criteria of effective hyperuniformity and stealthiness. When considering the system as a whole, which includes contributions from free particles, and effects of cluster size polydispersity, results are less conclusive. Probably larger samples might throw some light onto this problem, but they are beyond our current computational capabilities. One must bear in mind anyway that a complete study must incorporate the explicit influence of particle form factors, which play a fundamental role in determining the interaction with radiation probes. Using as starting point our glassy cluster fluid, for large colloidal particles ($\sigma\approx 1\sim 10\mu m$), attenuation would take place for radiation in the low frequency range of radio waves. The region of interest can be tuned resorting to mixtures of cluster-forming particles Chen et al. (2018) and/or controlling the size of the clusters. Future work will focus on mixtures where intracluster phase separation can enable the tuning of anisotropic cluster-cluster interactions and systems in which self-limitation is controlled by saturation of associative sites in order to reduce polydispersity.
Supplementary material
As supplementary information a graphic illustration of the aggregation process is included. In addition, velocity self-correlation functions are discussed an the net particle number fluctuation as a function of sampling radius is also introduced to complement the analysis in the Section V.
Acknowledgments
The authors acknowledge the support from the Agencia Estatal de Investigación and Fondo Europeo de Desarrollo Regional (FEDER) under grants No. PID2023-151751NB-I00 and PID2020-115722GB-C22. We also would like to acknowledge the Galicia Supercomputing Center (CESGA) for the access to their computer facilities.